40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 110 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  11992   Tue Feb 16 09:13:39 2016 SteveUpdateGeneralthere was a power outage & EQ

Pasadena reported the Sunday night event as a power transient caused by the trip of a power transmission line. This affected the entire city. Once the loss was detected, the system automatically switches to an alternate source. It was about one second for the system to recover.

 

2W Innolight, both Lightwaves at the ends, PSL HEPA, Ref Cavity and 3 AC units  turned on.

The 40m vacuum did not trip. It is vacuum normal.

The Jetstore computer is indicating power failer.

 

EQ  3.9 4.9 km (3.1 mi) NNW of Big Bear City, CA34.3027, -116.863, 3km Feb 16 2016 09:24:21 UTC 37524376

 

Attachment 1: 20160215_power_outage.png
20160215_power_outage.png
  11991   Mon Feb 15 13:09:33 2016 KojiUpdateGeneralSomething has gone wrong - was there a power outage?

Looks like that's the case. LIGO GC also sent an e-mail that there was a popwer glitch.

  11990   Mon Feb 15 12:28:03 2016 gautamUpdateGeneralSomething has gone wrong - was there a power outage?

I came into the 40m a few minutes ago, and noticed the following (approximately in this order):

  • The striptool plots projected onto the wall were gone, even though the projector seemed to be working fine
  • There was no light at all in the IFO 
  • There was an incessnt beeping noise coming from inside the lab.

To investigate further, I checked today's summary pages, and whatever caused this, happened around 730am today morning (approx 5 hours ago). I also saw that all the watchdogs were tripped, except MC3, BS and SRM. 

I then tracked down the beeping - I believe that it is coming from Megatron.(in fact, it is coming from the Jetstor..) 

I also found that the PSL is OFF, and the Marconi, though ON, has the display parameters set to values that I normally see when it is first turned ON (i.e. the carrier frequency is 1200MHz, the output is -140dBm etc - this is what led me to suspect that somehow the power connection was interrupted? As far as the workstation computers are concerned, I don't think ROSSA was affected, but pianosa is frozen and donatella is at the login screen. The CDS overview MEDM screen refuses to load correctly (though some of the other MEDM screens are working fine). I'm not entirely sure how to go about fixing all of this, so for now, I'm leaving the PSL off and I've shutdown the remaining watchdogs.

It just occurred to me to check the status of the vacuum - the MEDM screen seems to suggest everything is fine (see Attachment #1). I went down to the X end to do a quick check on the status of the turbo pumps and everything looks normal there...

Attachment 1: Vac_15Feb2016.png
Vac_15Feb2016.png
  11989   Fri Feb 12 19:07:52 2016 KojiUpdateGreen LockingLaser swap - green light recovered but no flashes in the arm

800e-6 / 0.225^2 = 0.016

=> 1.6%/W

I thought Kiwamu had roughtly 2%/W.

 

  11988   Fri Feb 12 17:05:40 2016 gautamUpdateGreen LockingLaser swap - green light recovered but no flashes in the arm

After carefully tweaking the mode-matching of the IR into the crystal and the four-axis translation stage on which the doubling oven is mounted, I managed to recover 800uW of green power going into the green Faraday. Considering we have ~225mW of IR power coming out of the IR faraday (and roughly that amount going into the SHG crystal), I'd say this is pretty consistent (if not slightly better) with a recent power budget I had made for the X end. The amount of green power we get out of the doubling crystal is very sensitive to the alignment of the crystal to the beam axis. I suspect we could improve the situation slightly if the mode-matching lenses were mounted on translational stages so we could tweak their position, but the current situation on the X endtable does not provide space for this. In any case, I'd say we are at least as good as we were before, and so this should be an adequate fix until the new end-table is installed (though I don't know why we aren't seeing the predicted SHG conversion efficiency of 3-4% as predicted by Kiwamu's calculations, we are getting more like .36% conversion efficiency)...

Because the alignment of the beam before the doubling oven had changed, I had to adjust the steering mirrors to make the green beam go into the green faraday. I had placed a couple of irides for the green beam as a reference of the old path into the arm, and I used these to adjust some of the green mirrors to center the green beam on these. However, I did not observe any flashes in the arm. I will check if we are still mode-matched to the arm, and if the lenses downstream of the doubling oven need to be moved....

  11987   Fri Feb 12 11:10:49 2016 SteveUpdateGreen LockingInnolight laser is 10 years old

It shipped out for repair evaluation.

Arrived to Hayward,CA   2016Feb16

 

Attachment 1: inno1W.jpg
inno1W.jpg
  11986   Thu Feb 11 14:28:50 2016 SteveUpdateTreasure091415 declared

   Beautifully Done

   Chirp

  what is next?

Atm 3, Ron Drever could not celebrate with us because of health issues.

 

Attachment 1: 091415declared.jpg
091415declared.jpg
Attachment 2: You_were_right!.jpg
You_were_right!.jpg
Attachment 3: P1080312.JPG
P1080312.JPG
  11985   Wed Feb 10 17:57:15 2016 gautamUpdateGreen LockingLaser swap - updates

After the discussion at the meeting, I decided to go ahead and open the top of the oven so that I could get a visual on where the crystal was located - this helped in the alignment, and I was able to get some green light out of the oven. I had to tweak the position of the Doubling oven a little (with the top open) in order to align the crystal to the beam axis. However - I was only able to get ~140uW of green light going into the Faraday. I had measured the power at various points along the beam path recently with the old setup. We used to have ~860uW of green going into the Faraday there. To see if I could improve the situation a little, I checked that the beam was reasonably centered on both apertures of the IR Faraday, and then removed the irides upstream of the doubling oven. These were preventing me from placing the lenses exactly as per the a la mode solution. Once the irides were removed, I moved the lenses to their optimal positions as best as I could with a tape measure to mark out distances. I then further tweaked the position of the doubling oven using the 4 axis stage, monitoring the green power while doing so. The best I could get was ~200uW. Perhaps the positions of the lenses need to be optimized further. I also checked the IR power before and after the IR Faraday - these numbers are ~260mW and ~230mW respectively (I maximized the transmitted power through the Faraday by rotating the HWP, the QWP that was in the beam path has now been removed as the Lightwave outputs linearly polarized light), and compare favourably to the numbers in the old setup. Doing a naive scaling accounting for the fact that we have less power going into the doubling crystal, I would expect ~700uW of green light coming out, so it looks like the mode matching into the doubling crystal is indeed sub-optimal. However, now that things are roughly aligned, I hope the optimization will go faster...

  11984   Tue Feb 9 19:15:36 2016 gautamUpdateGreen LockingLaser swap - updates

Some updates on the laser swap situation:

  1. Mode-matching calculation: 
  • I should have caught this earlier, but it was an oversight - the 35um waist that Andres used in his calculation is the waist size of the green beam. So I've been off by a factor of sqrt(2) all this while, and it works out that the desired waist size is indeed 50um, consistent with Kiwamu's elogs. Furthermore, as he has detailed in that elog, we actually want the free-space waist of the input beam to the doubling crystal to be ~6.7mm from the geometric center of the PPKPT crystal. 
  • I redid the calculation using these updated numbers. Attachment #1 shows the results (optimized for the X-waist, Y-profile plotted for comparison and to see what mode-matching efficiency we get). The way I've set up the code is for a la mode to rank the solutions in order of increasing sensitivity to the positions of the lenses. It turns out the least sensitive solution doesn't actually achieve the desired waist size of 50um - moreover, it requires us to change both lenses currently in the path. The next lease sensitive solution, however, achieves the desired waist (i.e. 100% theoretical mode-matching efficiency for the X mode) and only requires us to swap the 125mm lens we put in yesterday for a 150mm lens (and the positions of the lenses change slightly compared to what we had yesterday as well). The sensitivity in a la mode is parametrized by the amount of power remaining in the TEM00 mode while displacing one or more components. It turns out that this figure of merit is only ~1% smaller for the 2nd least sensitive solution compared to the first. So I've chosen to use that solution. Code used to calculate the mode matching is Attachment #2.
  • I've also plotted in Attachment #1 what the beam profile would have looked like before our modificatons last night, using the numbers from Zach's elog - as I have already mentioned in the previous elog, it suggests that the waist size would have been 39um, at a location 1.0821m in my coordinate system (desired position according to considerations in the previous 2 bullets is 1.0713). This seems to have been a sub-optimal configuration, but is also subject to errors I made in measuring the positions of the mirrors/lenses (I don't think I had 1cm resolution).

       2. Implementing the new solution:

  • I've switched out the 125mm efl lens for a 150mm efl lens from the same Thorlabs lens kit. I've also moved both the lenses to their new appropriate positions.
  • Unfortunately, I had put in some irides in the beampath before calculating this new (more appropriate solution). As a result, both the lenses are off from their optimal positions by a few mm because the irides get in the way. I guess we just have to live with this for now, and can adjust the positions of the lenses once we actually get some green light and are happy with all the other alignments...
  • As noted in the previous elog, I suspect that we saw no green light yesterday because we were missing the doubling crystal altogether (given that we have only a 1mm x 1mm area to aim for - the Faraday serves as a coarse constraint, though its aperture has ~25times this area!). I tried playing around with the two steering mirrors immediately after the NPRO to see if I could get some green light out, but have not been successful yet. I may make some further trials later in the evening/tomorrow...

As I check the manual of the Innolight (pg17) and the datasheet of the Lightwave, I wonder if the Quarter Wave Plate that was placed immediately after the Innolight laser head is even necessary now - I assume the purpose of the combination of QWP+HWP was to turn the elliptically polarized light from the Innolight into linearly polarized light before the Faraday. But the Lightwave already produces linearly polarized light. I will check out what is the configuration on the Y-end table...

 

Attachment 1: Modematch_X.pdf
Modematch_X.pdf
Attachment 2: XendModeMatch.m.zip
  11983   Tue Feb 9 11:49:47 2016 gautamUpdateGreen LockingLaser swap initiated

Steve pointed me to an old elog by Zach where he had measured the waist of the 1W Innolight NPRO. I ran a la mode with these parameters (and the original optics in their original positions prior to last night's activities), and the result is in reasonably good agreement (see Attachment #1) with my initial target waist of 35 um at the center of the doubling oven (which I presume coincides with the center of the doubling crystal). The small discrepancy could be due to errors in position measurement (which I did by eye with a tape measure) or because I did not consider the Faraday in the a la mode calculation. However, I wonder why this value of 35 um was chosen? In this elog, Kiwamu has determined the optimal waist size to be 50um at the center of the doubling crystal. Nevertheless, as per his calculations, the doubling efficiency should be non-zero (about 1% lower than the optimum conversion efficiency) at 35um or 70um, so we should be able to see some green light as long as we are in this fairly large range. So perhaps the fact that we aren't seeing any green light is down to sub-optimal alignment? I don't think there is a threshold power for SHG as such, its just that with lower input power we expect less green light - in any case, 200mW should be producing some green light... From what I could gather from a bunch of old elogs by Aidan, the Raicol PPKPT crystals have dimension 1mm x 1mm x 30mm (long axis along beam propagation), so there isn't a whole lot of room for error perpendicular to the direction of propagation... I wonder if it is possible, for the initial alignment, to have the top cover of the doubling oven open so that we can be sure we are hitting the crystal?

Attachment 1: Innolight_beamProfile.pdf
Innolight_beamProfile.pdf
  11982   Tue Feb 9 04:37:10 2016 ericqUpdateGreen LockingLaser swap initiated

[ericq, Gautam]

Tonight we embarked on the laser swap. In short, we have gotten ~210mW through the faraday doubler, but no green light is apparent. The laser outputs ~300mW, so it's not exactly a work of art, but I still expected some green. More work remains to be done...

Gautam took numerous photos of the table before anything was touched. One lens was swapped, as per Gautam's plan. The innolight laser and controller are on the work bench by the end table. The lightwave is on the table and on standby, and is not hooked up to the interlock mounted on the table frame, but instead one below the table directly next to the controller. The ETMX oplev laser is turned off. 

  11981   Mon Feb 8 15:36:37 2016 gautamUpdateGreen LockingAlternative mode-matching scheme

I looked in the optics cabinet to see what lenses we have available, and re-ran the mode-matching calculation to see if we could find a better solution - I'm attaching a plot for what looks like a good candidate (optimized mode-matching efficiency for the X mode is 100%, and for the Y mode, it is 97.98%), though it does involve switching "L1", which is currently a 175mm efl lens, for a 125mm efl lens. I've also indicated on the plot where the various other components are relative to the optimized positions of the lens, and it doesn't look like anything is stacked on top of each other. Also, the beam width throughout is well below 4.7mm, which is the maximum cited width the Faraday can handle, as per its datasheet. "L1" doesn't quite get the waist of the beam to coincide with the geometrical center of the Faraday, but I don't think this is requried? Also, I've optimized the mode matching using the measured X width of the beam (red curve in Attachment #1), and have overlaid the calculated Y width of the beam for the optimized position of the lenses (red curve in Attachment #1). The target waist was 35um at the center of the doubling oven, which the X profile achieves, but the Y profile has a width of 32 um at the same point.

In all the calculations, I've not accounted for possible effects of the HWPs and the Faraday on the beam profile....

Attachment 1: Modematch_alternative.pdf
Modematch_alternative.pdf
  11980   Mon Feb 8 09:37:07 2016 SteveUpdateSUS EQ 3.9m

No sign of damage

 

Attachment 1: EQ3.9mSimmierCa.png
EQ3.9mSimmierCa.png
  11979   Fri Feb 5 16:50:24 2016 gautamUpdateGreen LockingFirst pass at mode-matching

I've done a first pass at trying to arrive at a mode-matching solution for the X-end table once we swtich the lasers out. For this rough calculation, I used a la mode to match my seed beam (with z = 0 being defined as the shutter housing on the current position of the Innolight laser head, and the waist of the beam from the NPRO being taken as the square-root of the X and Y waists as calculated here), to a target beam which has a waist of 35um at the center of the doubling oven (a number I got from this elog). I also ignored the optical path length changes introduced by the 3 half-wave plates between the NPRO and the doubling oven, and also the Faraday isolator. The best a la mode was able to give me, with the only degrees of freedom being the position of the two lenses, was a waist of 41um at the doubling oven. I suppose this number will change once we take into account the effects of the HWPs and the Faraday. Moreover, the optimized solution involves the first lens after the NPRO, L1, being rather close to the second steering mirror, SM2 (see labels in Attachment #2, in cyan), but I believe this arrangement is possible without clipping the beam. Moreover, we have a little room to play with as far as the absolute physical position of the z=0 coordinate is - i.e. the Lightwave NPRO head can be moved ~2cm forward relative to where the Innolight laser head is presently, giving a slightly better match to the target waist (see attachment #3). I will check the lenses we have available at the 40m to see if a more optimal solution can be found, but I'm not sure how much we want to be changing optics considering all this is going to have to be re-done for the new end table... Mode-matching code in Attachment #4...

Attachment 1: Modematch_AUXx.pdf
Modematch_AUXx.pdf
Attachment 2: NewSetUp.png
NewSetUp.png
Attachment 3: Modematch_AUXx_2.pdf
Modematch_AUXx_2.pdf
Attachment 4: XendModeMatch.m.zip
  11978   Fri Feb 5 15:02:13 2016 gautamUpdateGreen LockingX-end NE cable

[Steve, gautam]

Steve thinks that the X-end Innolight does not come with the noise-eater option (it is an add-on and not a standard feature, and the purchase order for the PSL Innolight explicitly mentions that it comes with the NE option, but the X-end Innolight has no such remarks), which would explain why there is no difference with the noise eater ON/OFF. During earlier investigations however, I had found that there was a cable labelled "Noise-Eater" connected to one of the Modulation Inputs on the rear of the Innolight controller. Today, we traced this down. The modulation input on the rear says "Current Laser Diode 0.1A/V". To this input, a Tee is connected, one end of which is terminated with a 50ohm terminator. The other end of the Tee is connected to a BNC cable labelled "Nosie-Eater", which we traced all the way to the PSL table, where it is just hanging (also labelled "X end green noise eater"), unterminated, at the southeast corner of the PSL table. It is unlikely that this is of any consequence given the indicated coefficient of 0.1A/V, but could this somehow be introducing some junk into the laser diode current which is then showing up as intensity fluctuations in the output? Unfortunately, during the PLL measurements, I did not think to disconnect this BNC and take a spectrum. It would also seem that the noise-eater feedback to the laser diode current is implemented internally, and not via this external modulation input jack (the PSL, which I believe has the noise-eater enabled, has nothing connected to this rear input)...

 

  11977   Fri Feb 5 00:23:01 2016 gautamUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

I've repeated the measurement for the x-direction and also did the y-direction, taking into account Koji's suggestion of keeping the power meter as close as possible to the knife edge. Attachment #1 shows a picture of the setup used. Because an ND filter is required to use this particular power meter, the geometrical constraints mean that the closest the power meter can be to the knife edge is ~3cm. I think this is okay. 

The result from the re-measured X-scan (Attachments #2 and #4) is consistent with the result from yesterday. Unfortunately, in the y-direction (Attachments #3 and #4), I don't seem to have captured much of the 'curved' part of the profile, even though I've started from pretty much adjacent to the HWP. Nevertheless, the fits look reasonable, and I think I've captured sufficient number of datapoints to have confidence in these fits - although for the Y-scan, the error in the waist position is large. The ellipticity as measured using this method is also significantly smaller than what the CCD beam profiler was telling us. 

If we are happy with this measurement, I can go ahead and work on seeing if we can arrive at a minimally invasive mode-matching solution for the X-end table once we switch the lasers out...

 

Attachment 1: Beamscan_setup.pdf
Beamscan_setup.pdf
Attachment 2: Beamscan_x.pdf
Beamscan_x.pdf
Attachment 3: Beamscan_y.pdf
Beamscan_y.pdf
Attachment 4: Zscan.pdf
Zscan.pdf
  11976   Thu Feb 4 10:19:05 2016 SteveUpdatesafetyfire marshal inspection

Pasadena fire marshal inspected the lab today. No violation was found.

  11975   Thu Feb 4 10:14:52 2016 SteveUpdateSUS earthquake 3.5M

 

 

Attachment 1: EQ3.5m_Simmier.Ca.png
EQ3.5m_Simmier.Ca.png
  11974   Thu Feb 4 09:16:46 2016 KojiUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

I don't think the discrepancy is a serious issue as long as the mode is clean. The mode is determined by the NPRO crystal and is hard to change by anything except for the thermal lensing in the crystal.

And I never succeeded to reproduce the mode listed in the manual.

One thing you'd better to take care is that clipping of the beam produces diffraction. The diffracted beam spreads faster than the nominal TEM00 mode. Therefore the power meter should to be placed right after the razor blade. i.e. As you move the longitudinal position of the razor blade, you need to move the power meter.

  11973   Wed Feb 3 23:23:47 2016 gautamUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

As Koji pointed out in the previous elog, the CCD beam profiler was ill suited for this measurement. Nevertheless, to get a rough idea of the beam profile, I made a few rearrangements to my earlier setup:

  • Kept the HWP at the same place it was, as this is roughly the configuration that is going to be used at the endtable anyways. It was ~7cm from the shutter housing on the laser head (unfortunately, I neglected to take a picture).
  • Moved the PBS downstream till it was ~40 cm away (so as to minimize the thermal lensing effect from the ~300mW beam) from the laser head. Rotated the HWP till I got about 6mW of transmitted power (dumped the rest into a black hole)
  • Installed a 95% reflecting BS to further attenuate the power to a level suitable for the CCD (dumped the reflected part onto a razor beam dump)
  • Installed the CCD beam profiler and captured an image, at ~60cm from laser head. In this configuration, I was able to get a clean image capture without the CCD saturating. Unfortunately, I could not transfer it off the laptop used to operate the beam profiler, I will upload a screen capture tomorrow once I get it. Anyways, the main observation was that the beam appeared quite elliptical (ellipticity ~0.6). It was also not clear to what extent thermal lensing at the PBS/BS was afftecting this measurement.

Following Koji's suggestion, I decided to do a knife-edge measurement as well. The measurement configuration was similar to the one described above, except the PBS/BS were removed, and a 1.0 neutral density filter was was installed ~80cm from the laser head (here the ~300 mW beam was >2mm in diameter, as judged by eye). I used the Ophir power meter, which was why I had to install an ND filter as it is rated for 100mW max power. I will put a picture up tomorrow. Thermal lensing shouldn't be of much consequence here, as we just need the whole beam to fall onto the power meter active area (verified by eye), and only the relative change in power levels as the knife edge cuts the beam matters. I took the cross-sectional profile of the beam by translating the knife in the x-direction (i.e. cut the beam "left to right" ).

Attachments 1 and 2 are the results from todays measurements. It remains to repeat by cutting the beam along the y direction, and see what ellipticity (if any) shows up. I also found some "nominal" numbers in page 4 of the Lightwave datasheet - it tells us to expect a waist 5cm from the shutter housing, with horizontal and vertical 1/e^2 diameters of 0.5mm and 0.38mm respectively. My measurement suggests a horizontal diameter of ~0.25mm (half the "nominal" value?!), and the waist location to be 8.22cm from the shutter housing. I wonder if this discrepancy is a red flag? Could it be due to the HWP? I'm reasonably sure of my calculations, and the fits have come out pretty nicely as well...

Attachment 1: KnifeEdgeScans_x.pdf
KnifeEdgeScans_x.pdf
Attachment 2: Beamscan_x.pdf
Beamscan_x.pdf
  11972   Wed Feb 3 08:39:17 2016 SteveUpdateCDSblank daily summary pages

Daily summary pages are blank today. Yesterday is ok

  11971   Tue Feb 2 18:54:02 2016 KojiUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

jiIn fact, it is one of the most difficult type mode profiling to measure a beam directly out from a laser source.

If you reduce the power by ADJ, this significantly changes the output mode as the pumping power varies temperature gradient of the laser crystal and thus thermal lensing in it. I'd recommend you to keep the nominal power.

If you use a PBS for power reduction, you should increase the transmission ~x10 from the minimum so that you are not dominated by possible junk polarization.

Any transmissive BK7 components where the beam is small can cause thermal lensing. In order to avoid this issue, I usually use two noncoated (or one AR coated) optical windows made of UV fused silica to pick off the beam. Once the beam power is reduced I suppose it is OK to use an additional ND filter in front of the CCD.

Another more reliable method is an old-good knife edge measurement.

  11970   Tue Feb 2 18:35:47 2016 gautamUpdateGreen LockingLightwave NPRO moved from PSL table to SP table

I've moved the following components that was a part of Koji's setup from the PSL table to the SP table so that I may measure the beam profile of the beam from the spare Lightwave NPRO and work on a mode-matching solution for the X-end.

  • Lightwave laser
  • Lightwave controller
  • Interlock switch (Newport)
  • HWP and PBS

I did some preliminary characterization of the beam from the Lightwave - in the power controlled mode, setting the "ADJ" parameter to 0 (which is the state recommended in the manual) gives an output power of ~240mW. I used the HWP and PBS to dump most of this into a "Black Hole" beam dump, but I was still getting about 300uW of power after this. This was saturating the CCD in the beam profiler (even though 300uW for a beam of ~1mm should be well within the recommended operating limits as per its manual - maybe the ND filter on the camera isn't really ND4.0), and so I further reduced the "ADJ" parameter on the laser controller to -20, such that I had no saturation of the CCD. I will try and take some data later today. The laser is presently in "Standby" mode, and the SP table is fully covered again.

  11969   Mon Feb 1 18:11:25 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

Before distrubing the beat setup with the spare Lightwave laser, I wanted to see if I could resolve the apparent difference in behaviour between the measured free running noise of the spare Lightwave laser and my earlier measurements with the existing X and Y end lasers above ~5kHz. So I redid the measurement, but this time, on Eric's suggestion, while taking spectra on the SR785, I was careful to maintain the same "CH1 input range" while measuring the control signal spectrum and the measurement noise spectra. The level used was -20dBvpk. I think the measured spectrum shape now makes sense - above ~4kHz, the SR560 noise means that the SNR is poor and so we can only trust the spectra up to this value (the spectra for the end lasers are from earlier measurements where I did not take care to keep the input range constant). Anyways, I think the conclusion is that the spare Lightwave seems to have a free-running frequency noise that is approximately a factor of 3 worse than the Lightwave laser at the Y-end, though this may be because I didn't take the measurement at the optimal operating conditions (diode current, power etc). But I guess this is tolerable and that we can go ahead with the planned swapping out of the existing Innolight at the X-end with this laser. 

I will now move the Lightwave laser off the PSL table onto the SP table where I will do some beam characterization and see if I can come up with a satisfactory mode-matching solution for the swap. I've borrowed a beam profiler from the TCN lab for this purpose.

Attachment 1: Free_running_frequency_noise_comparison.pdf
Free_running_frequency_noise_comparison.pdf
  11968   Mon Feb 1 15:43:18 2016 KojiUpdateGreen LockingInnolight laser is 10 years old

This is the same one as what you got from Steve. But you can find full pages.

https://wiki-40m.ligo.caltech.edu/PSL/NPRO

  11967   Mon Feb 1 15:16:28 2016 gautamUpdateGreen LockingInnolight laser is 10 years old

The Innolight laser control unit has a 25 pin D-sub connector on the rear which is meant to serve as a diagnostics aid, and the voltages at the various pins should tell us the state of various things, like the diode power monitor, laser crystal TEC error temperature, NE status etc etc. Unfortunately, I am unable to locate a manual for this laser (online or physical copy in the filing cabinets), so the only thing I have to go on is a photocopied page that Steve had obtained sometime ago from the manual for the 2W NPRO. According to that, Pin 1 is "Diode laser 1, power monitor, 1V/W". The voltage I measured (with one of the 25 pin breakout boards and a DMM) is 1.038V. I didn't see any fast fluctuations in this value either. It may be that the coefficient indicating "normal" state of operation is different for the 1W model than the 2W model, but this measurement suggests the condition of the diode is alright after all?

I also measured the voltage at Pin 12, which is described in the manual as "Noise Eater, monitor". This value was fluctuating between ~20mV and ~40mV. Toggling the NE switch on the front of the control unit between ON and OFF did not change this behaviour. The one page of the manual that we have, however, doesnt provide any illumination on how we are supposed to interpret the voltage measured at this pin...

  11966   Mon Feb 1 14:08:06 2016 SteveUpdatePEMroof condtion is good

It rained hard yesterday. We have not had this strong downpoor for years. We got 0.7" and the roof did not leak.

  11965   Mon Feb 1 09:16:32 2016 SteveUpdatePEMRat got cut

We got it! Traps are removed.

 

Attachment 1: bingo.jpg
bingo.jpg
Attachment 2: bingo3.png
bingo3.png
  11964   Sat Jan 30 09:56:24 2016 KojiUpdateGreen LockingInnolight laser is 10 years old

It is strange that there is no difference between with and without NE, isn't it?

  11963   Sat Jan 30 00:12:22 2016 gautamUpdateGreen LockingInnolight laser is 10 years old

 

Quote:

I don't think there's any evidence that the noise eater is bad. That would change the behavior of the relaxation oscillation which is at 1 MHz ?

While I was investigating the AM/PM ratio of the Innolight, I found that there was a pronounced peak in the RIN at ~400kHz, which did not change despite toggling the noise eater switch on the front panel (see plot attached). The plot in the manual suggests the relaxation oscillations should be around 600kHz, but given that the laser power has dropped by a factor of ~3, I think it's reasonable that the relaxation oscillations are now at ~400kHz? 

Attachment 1: RIN_comparison.pdf
RIN_comparison.pdf
  11962   Fri Jan 29 16:55:27 2016 ranaUpdateGreen LockingInnolight laser is 10 years old

I don't think there's any evidence that the noise eater is bad. That would change the behavior of the relaxation oscillation which is at 1 MHz ?

  11961   Fri Jan 29 14:43:47 2016 SteveUpdateGreen LockingInnolight laser is 10 years old
Quote:

After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I

Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010

It's diodes should be replaced, based on it's age and performance.

RIN and noise eater bad. I will get a quote on this job.

The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956

Attachment 1: inno1W.pdf
inno1W.pdf
  11960   Fri Jan 29 09:59:23 2016 SteveUpdateVACvacuum control upgrade

Rana stated yesterday that there will be a vacuum control update in the close future. Witnesses : Rich, Chris and Dave

Can you give me this in writing?

 

Attachment 1: 1100days.png
1100days.png
  11959   Fri Jan 29 08:18:13 2016 SteveUpdatePEMPEM

 

Quote:

Air condition maintenance is happening. It should be done by 10am

 

Attachment 1: PEM.png
PEM.png
  11958   Thu Jan 28 19:10:16 2016 gautamUpdateLSCStatus of the green PDH circuits

 

Quote:

We will update the X circuit DCC page with an accurate schematic and photo. 

I've uploaded reasonably high-resolution photographs of the uPDH box for the X-end and Y-end on their respective wiki pages. I've uploaded two photos for each box, one of the circuit board (I checked that these photos are clear enough that we can zoom in and read off component values if necessary), and one of the box with the peripherals not integrated into the circuit board (i.e. the minicircuits mixer ZAD-8+ and the little Pomona box that is an LP filter for the output from the mixer). Since I pulled the boxes out, I thought it might not be a bad idea to measure the TFs of these Pomona boxes and make sure nothing weird is going on, I'll put up some plots later. 

Rana and I discussed some things to look at earlier today:

  • Check the voltages at test-points 1,7 and 9, and make sure they are as expected
  • Change the gain of the pre-amp from x2 to x4 - as Eric mentioned in the previous elog, these had already been swapped out. Right now a 600 ohm and 200 ohm resistor pair are being used, so the preamp gain is x4, which should be okay as per the datasheet of the AD8336 VGA amplifier (although the "recommended" resistor pairing is 301 ohm and 100 ohm, but I don't think this is critical?
  • Track down the reason for the difference in Gain settings at the X and Y ends - typically, the X-end PDH box "Gain" knob is set to 10, while that for the Y-end is set to ~4. The fact that the PZT actuator gain for the Y-end is ~5 times larger than for the X-end doesn't seem to account for all of this. As per the attached plot, the difference in gain between the ends is ~35 dB, which is a factor of 50!

I also did a quick check of the behaviour of the Servo Gain potentiometer by checking the resistance at various positions of the knob - we had suspected that the potentiometer may be logarithmic, but I found that it was in fact linear. I'll put up a plot of the gain as a function of the Servo Gain knob position soon,(plot added) along with results from the other checks.

While disassembling the setup at the X-end to get the PDH box out, I noticed that the signal from the LO is going to the mixer through a Pomona box (no such Pomona box is used at the Y-end). I opened it up and found that it contains just a pair of capacitors in parallel, so it's a phase shifter?. The LO signal also goes through an attenuator. The mixer in both boxes is a ZAD-8+, so why is this part of the setup different?

Both PDH boxes are not hooked up at the moment, I will restore the setups at both ends after running a few more checks on the boxes...

Attachment 1: Servo_gain_calibration.pdf
Servo_gain_calibration.pdf
  11957   Thu Jan 28 14:54:49 2016 ericqUpdateLSCStatus of the green PDH circuits

Yesterday, I uploaded some EAGLE schematic files and a LISO source file for the green PDH servo electronics to the 40m LISO git repository. In doing so, I realized that the DCC document for the X box (D1400293) was not updated at the end of the electronics work we did in Aug/Sep 2014. This is entirely my fault. 

The Y box document (D1400294) is currently accurate. 

The missing information is that, as I posted In ELOG 10457, I ended up destroying our original X box, and replaced it with a spare from the ATF. It was restuffed to match the Y end box pretty much exactly. We will update the X circuit DCC page with an accurate schematic and photo. 

Gautam tells me that he and Rana were looking at the outdated schematic and thinking about improvements, but at least some of this was already done back in 2014 (specifically, the resistors used to specify the AD8336 preamp gain were changed).

  11956   Thu Jan 28 00:29:30 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

Summary of the work done today:

Alignment and other work on PSL table

As mentioned in a previous elog, the beatnote amplitude I obtained was tiny - so I checked the alignment of the two beams onto the PD. I did this as follows:

  • Checked the alignment of the two beams on the recombination BS. Moved the steering mirror for the PSL beam until the two were aligned, as verified by eye using an IR card
  • Turn the steering mirror just before the fast focusing lens and thorlabs PD (kept the fork fixed, just loosened the screw on the post to do this) such that the far-field alignment of the two beams could be checked. I used the BS to tweak this alignment as necessary
  • Iterate the previous two steps till I was happy with the alignment
  • Return steering mirror before the PD to its original position, tweak alignment until DC level on the PD was maximized (as verified using an oscilloscope) 
  • Adjust the HWP just after the lightwave laser such that the power arriving at the PD from the PSL beam and the lightwave beam were approximately equal - verified by blocking each beam and checking change in the DC level

After doing all of this, I found a beatnote at ~-10dBm at a temperature of 45.3002 degrees on the Lightwave. The DC level was ~8V (~4V contribution from each beam). 

PLL and frequency nosie measurements:

Pretty much the same procedure as that described in this elog was followed for setting up the PLL and taking the measurements, except that this time, I used the two SR560s in a better way to measure the open loop TF of the PLL. This measurement suggested a UGF of ~ 10kHz, which seems reasonable to me. I turned the 11MHz marconi off because some extra peaks were showing up in the beat signal spectrum. I judged that the beatnote was not large enough to require the use of an attenuator between the PD and the mixer. I was able to lock the PLL easily enough, and I've attached spectra of the control signal (both uncalibrated and calibrated). To calibrate the spectrum, I did a quick check to determine the actuator gain of the spare Lightwave laser, by sweeping the fast PZT with a low frequency (0.5Hz) 1Vpp sine wave, and looking at the peak in the beat signal spectrum move on the network analyzer. This admittedly rough calibration suggests that the coefficient is ~5MHz/V, consistent with the other Lightwave. Eric suggested a more accurate way to do this would be to match up spectra taken using this method and by locking the PLL by actuating on the FM input of the Marconi - I didn't try this, but given the relatively large low-frequency drifts of the beatnote that I was seeing, and that the control signal was regularly hitting ~2V (i.e shifting the frequency by ~10MHz), I don't think this is viable with a low MHz/V coefficient on the Marconi, which we found is desirable as described here

Bottom line:

The spare Lightwave frequency noise seems comparable to the other two measurements (see attachment #2). If anything, it is a factor of a few worse, though this could be due to an error in the calibration? I'm also not sure why the shapes of the spectra from today's measurement differ qualitatively from those in elog 11929 above ~7kHz. 

 

Some random notes:

  • Do we want to do an AM/PM characterization of the spare Lightwave laser as well? It might be easier to do the PM measurement while we have this measurement setup working
  • Yesterday, I noticed some peaks in the spectrum of the PD output while only the PSL beam was incident on it, at ~35MHz and ~70 MHz. They were pretty small (~-50dBm), but still clearly discernible over the analyzer noise floor. It is unclear to me what the source of these peaks are.
Attachment 1: PLL_OLG.pdf
PLL_OLG.pdf
Attachment 2: Freq_noise_comparison.pdf
Freq_noise_comparison.pdf
  11955   Wed Jan 27 23:14:25 2016 gautamUpdatePEMETMX floor vs table noise
Quote:

I didn't really appreciate this measurement until just now. IF you can save the DTT .xml file with all the traces in it (i.e. NOT just the plots), we should save this data for comparison plotting later. Perhaps Gautam can post the gzipped xml file for you into the log.

The accelerometers don't read any real noise below ~3 Hz, so we can't judge the difference down low, but this seems like a good measurement in the 5 - 100 Hz band.yes

Unfortunately I had closed all the DTT windows that Steve had used for the earlier plots. So I took the spectra again - there may be minor differences given that this measurement was taken at ~11pm at night. Anyways, plots and the xml data file are attached.  

 
Attachment 1: X_end_ACC_Data.png
X_end_ACC_Data.png
Attachment 2: X_End_ACC_Data.xml.tar.gz
  11954   Wed Jan 27 20:51:00 2016 ranaUpdatePEMETMX floor vs table noise

I didn't really appreciate this measurement until just now. IF you can save the DTT .xml file with all the traces in it (i.e. NOT just the plots), we should save this data for comparison plotting later. Perhaps Gautam can post the gzipped xml file for you into the log.

The accelerometers don't read any real noise below ~3 Hz, so we can't judge the difference down low, but this seems like a good measurement in the 5 - 100 Hz band.yes

  11953   Wed Jan 27 18:19:45 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...

  11952   Wed Jan 27 08:17:38 2016 SteveUpdatePEMETMX floor vs table noise

 

Quote:

Objective:  measure the noise floor on the optical table and the floor so we can decide if the table needs better anchoring before swapping in

                   the larger optical table

The accelerometrs labeled as MC1 ( just north east  of IOO chamber floor ) and MC2 ( north east leg of MC2 table floor ) were moved:

MC1 to the floor at the north west leg of optical table.

MC2 is in the north east corner of the optical table

Atm2 was taken after table leg bolts were tighed at 40 ft/lb

The spectrum looks similar to ETMY       except  the Z direction 

ETMY .

Conclusion: up to 20 Hz this set up is good.

Attachment 1: ETMX_X_floor_vs_table.png
ETMX_X_floor_vs_table.png
Attachment 2: ETMX_Y_floor_vs_table.png
ETMX_Y_floor_vs_table.png
Attachment 3: ETMX_Z_floor_vs_table.png
ETMX_Z_floor_vs_table.png
Attachment 4: MC1@ETMX.jpg
MC1@ETMX.jpg
Attachment 5: MC2@1Winno.jpg
MC2@1Winno.jpg
Attachment 6: MC2@ETMXtable.jpg
MC2@ETMXtable.jpg
  11951   Tue Jan 26 17:50:22 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

I attempted to measure the frequency noise of the extra Lightwave NPRO we have that is currently sitting on the PSL table. I did the following:

  1. Turn the Lightwave NPRO back on.
  2. Disable MC autolocker and close the PSL shutter.
  3. Checked the alignment of the pick off from the PSL beam and the beam from the Lightwave NPRO onto the PDA10CF. These seemed okay, and I didn't really have to tweak any of the steering optics. I was getting a DC signal level of ~7V (the PD should drive a 1Mohm load up to 10V so it wasn't saturated).
  4. Swept the crystal temperature on the Lightwave using the dial on the front panel of the controller. I found beatnotes at 48.1831 degrees and 45.3002 degrees. However, the amplitude of the beatnote was pretty small (approx. -40dBm on the Agilent NA). I tried playing around with the beam alignment and laser power on the Lightwave NPRO to see if I could increase the beatnote amplitude, but was unsuccessful - turning up the laser power (from the nominal level of 55mW as per the front panel display) caused the PD to saturate at 10V, while as far as I could tell, the alignment of the two beams onto the PD is reasonably good. This seems inconsistent with the numbers Koji has reported in this elog, where he was able to get a beatnote of ~1Vpp for a DC of 2.5 V. 
  5. I tried locking the PLL (in roughly the same configuration as reported here) with this small amplitude beatnote but was unsuccessful. 

I've turned the Lightwave NPRO back to standby for now, in anticipation of further trials later today. I've also restored the IMC. 

  11950   Tue Jan 26 15:07:47 2016 SteveUpdatePEMaccelerometers moved to ETMX

Objective:  measure the noise floor on the optical table and the floor so we can decide if the table needs better anchoring before swapping in

                   the larger optical table

The accelerometrs labeled as MC1 ( just north east  of IOO chamber floor ) and MC2 ( north east leg of MC2 table floor ) were moved:

MC1 to the floor at the north west leg of optical table.

MC2 is in the north east corner of the optical table

Atm2 was taken after table leg bolts were tighed at 40 ft/lb

The spectrum looks similar to ETMY     (Guralp for below 3 Hz )  except  the Z direction 

ETMY  .

Attachment 1: ETMX_floor_vs_optab.png
ETMX_floor_vs_optab.png
Attachment 2: ETMX_floor_vs_optabtop.png
ETMX_floor_vs_optabtop.png
  11949   Mon Jan 25 11:06:08 2016 SteveUpdatePEMnoisy fan belt replaced

The noisy fan belt on the roof of the control room was finally replaced on Friday, Jan 22 2016

Quote:

Air condition maintenance is happening. It should be done by 10am

 

  11948   Mon Jan 25 08:36:46 2016 SteveUpdateSUSPRM damping restored

PRM suspension damping restored after 4.1 Mag Ludlow earthquake.

 

Attachment 1: EQ4.1mLudlow.png
EQ4.1mLudlow.png
Attachment 2: EQ7.1mAlaska_4.1mLudlow.png
EQ7.1mAlaska_4.1mLudlow.png
Attachment 3: 7.1_4.1.png
7.1_4.1.png
  11947   Fri Jan 22 18:46:03 2016 ranaUpdateGreen LockingAUX-X AM/PM investigations

The PDA photodetectors are DC coupled, so you cannot use them to go directly into the analyzer. Must use the DC block so that you can reduce the input attenuation on the B channel and then lower the drive amplitude.

Good policy for TF measurements: drive as softly as you can and still measure in a reasonable amount of time, but no softer than that.

  11946   Fri Jan 22 17:22:06 2016 gautamUpdateGreen LockingAUX-X AM/PM investigations

There were a number of directories in /users/OLD/mott/PZT/2NPRO, I've used the data in Innolight_AM_New. Also, I am unsure as to what their "calibration" factor is to convert the measured data into RIN, so I've just used a value of 0.8, with which I got the plot to match up as close as possible to the plot in this elog. I also redid the measurement today, given that the laser parameters have changed. The main difference was that I used an excitation amplitude of +15dBm, and an "IF Bandwidth" of 30Hz in the parameter files for making these measurements, which I chose to match the parameters Mott used. There does seem to be a shift in some of the features, but the <100kHz area seems similar to the old measurement now. 

Having put the PD back in, I also took measurements of the RIN with the input to the laser PZT terminated. There is no difference with the Noise Eater On or OFF! 

Quote:
Quote:

Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 

It looks like some of the features may have shifted in frequency. The previous measurement results can be found in /users/OLD/mott/PZT/2NPRO, can you plot the two AM measurements together?

 

Attachment 1: AM_response.pdf
AM_response.pdf
Attachment 2: NE_investigations.pdf
NE_investigations.pdf
  11945   Fri Jan 22 13:33:37 2016 ericqUpdateGreen LockingAUX-X AM/PM investigations
Quote:

Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 

It looks like some of the features may have shifted in frequency. The previous measurement results can be found in /users/OLD/mott/PZT/2NPRO, can you plot the two AM measurements together?

  11944   Fri Jan 22 11:33:20 2016 gautamUpdateGreen LockingAUX-X AM/PM investigations

I was trying to characterize the AM/PM response of the X end laser. I tried to measure the AM response first, as follows:

  • I used the Thorlabs PDA 55, whose datasheet says it has 10MHz bandwidth - I chose it because it has a larger active area than the PDA 255, but has sufficient bandwidth for this measurement. 
  • My earlier measurement suggested the IR power coming out of the laser is ~300mW. As per the datasheet of the PDA 55, I expect its output to be (1.5 x 10^4 V/A) * (~0.25 A/W) ~ 4000 V/W => I expect the PD output (driving the 50ohm input of the Agilent NA) to saturate at ~1.3mW. So I decided to use a (non-absorptive) ND 3.0 filter in front of the PD (i.e. incident power on the PD ~0.3 mW).
  • I measured the AM response (inputA/inputR) by using the RF output from the Agilent analyzer (divided using a mini-circuits splitter half to input R and half to the laser PZT), and the PD output to input A. I set the power of the RF output on the analyzer to 0 dBm. 
  • Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 
  • I also took a measurement of the RIN with no drive to the laser PZT (terminated with a 50ohm terminator) - see Attachment #2. Qualitatively, this looks like the "free-running" RIN curve on the Innolight datasheet (see Attachment #3, the peak seems slightly shifted to the left though), even though the Noise Eater switch on the laser controller front panel is set to "ON". I neglected taking a spectrum with it OFF, I will update this elog once I do (actually I guess I have to take both spectra again as the laser diode and crystal temperatures have since been changed - this data was taken at T_diode = 28.5deg, I_diode = 1.90A, and T_crystal = 47.5 deg). But does this point to something being broken?
  • I was unable to lock the PLL yesterday to measure the PM response, I will try again today.
Attachment 1: AUX_X_AM.pdf
AUX_X_AM.pdf
Attachment 2: AUX_X_RIN.pdf
AUX_X_RIN.pdf
Attachment 3: NE_Mephisto.png
NE_Mephisto.png
  11943   Fri Jan 22 01:56:01 2016 ericqUpdateLSCAudio ALSX

I hooked up the ALSX DFD output to the fibox, and used the adjustable delay line to set the phase properly. I recorded the noise on pianosa, and have attached it. Of course, this doesn't really capture the low frequency behavior. 

Unrelated to this: I found the MC WFS turned off, and the loops ran away when turning them on. I tweaked the alignment, and reset the WFS offsets. Seems stable for now. 

Attachment 1: ALSX.wav
ELOG V3.1.3-